OpenStudy (anonymous):

-4x+2<-1(-1x-1)+2 i got (-infty, -1/5] but it says its not correct

OpenStudy (anonymous):

\[-4x+2<x+1+2\] is a start

OpenStudy (anonymous):

i worked it out and got x<-1/5 is that right?

OpenStudy (anonymous):

\[-4x+2<x+3\] \[2<5x+3\] \[-1<5x\] \[-\frac{1}{5}<x\] which says x is GREATER than -1/5

OpenStudy (anonymous):

so as an interval it should be \[(-\frac{1}{5},\infty)\]

OpenStudy (anonymous):

oh i see that now, thanks